### v1

# thanks: These two authors contribute equally in this workthanks: These two authors contribute equally in this work

Abstract
Since the very recent discovery of unconventional superconductivity in twisted WSe2 homobilayers at filling $\nu=-1$, considerable interests arise in revealing its mechanism. In this paper, we developed a three-band tight-binding model with non-trivial band topology by direct Wannierization of the low-energy continuum model. Incorporating both onsite Hubbard repulsion and next-nearest-neighbor attraction, we then performed a mean-field analysis of the microscopic model and obtained a phase diagram qualitatively consistent with the experiment results. For zero or weak displacement field, the ground state is a Chern number $C=\pm 2$ topological superconductor in the Altland-Zirnbauer A-class (breaking time-reversal but preserving total $S_{z}$ symmetry) with inter-valley pairing dominant in $d_{xy}\mp id_{x^{2}-y^{2}}$-wave (mixing with a subdominant $p_{x}\pm ip_{y}$-wave) component. For a relatively strong displacement field, the ground state is a correlated insulator with $120^{\circ}$ antiferromagnetic order. Our results provide new insights into the nature of the twisted WSe2 systems and suggest the need for further theoretical and experimental explorations.
Introduction.—In recent years, twisted van der Waals moiré superlattices garnered significant interests following the groundbreaking discovery of unconventional superconductivity (SC) [1, 2, 3] in magic-angle twisted bilayer graphene [4]. Beyond graphene-based systems, twisted transition metal dichalcogenides (TMDs) bilayers [5, 6] are regarded as a promising alternative platform for investigating many-body physics. Due to their high tunability, twisted TMDs can host a wide variety of exotic phases, including correlated insulators [7, 8, 9, 10], integer and fractional quantum anomalous Hall states [11, 12, 13, 14], and also integer and fractional quantum spin Hall states [15]. However, the experimental realization of SC in twisted TMD systems has remained elusive [9], though many theoretical works have suggested that SC should develops upon doping the correlated insulators [16, 17, 18, 19, 20, 21, 22, 23]. Very recently, two independent studies have reported the discovery of robust SC phases in twisted WSe2 (tWSe2) homobilayers [24, 25]. The phenomenology of these SC phases differ significantly from that would arise in doped Mott insulators, which call for further theoretical investigations, particularly regarding the possible pairing mechanism and topological properties of SC, as well as the nature of adjacent correlated insulators.
Previous theoretical studies [9, 26, 27, 28, 16, 17, 19, 20, 21, 18, 22, 23] of tWSe2 homobilayers are largely based on the tight-binding description of the moiré Hubbard model with spin-dependent hopping phase tuned by the displacement field. This approach provides valuable real-space insights essential for understanding the correlated physics. However, such a simple one-band model cannot capture the potential nontrivial band topology [6, 29] and only applies to limited twist angles [9]. Thus, a direct Wannierization [29, 30, 31, 32, 33, 34] of the standard continuum model [6] with a proper set of low-energy bands appears to be a more suitable approach to start with.
In this paper, we focus mainly on the experimental results in Ref. [24], where the SC phase is observed in a $3.65^{\circ}$ tWSe2 device at integer filling factor $\nu=-1$ under small displacement field, along with a correlated insulator phase at larger displacement field. Using the continuum model parameters provided in Ref. [29], we first construct a three-band tight-binding model for $3.65^{\circ}$ tWSe2 through direct Wannierization [29, 30, 31, 32, 33, 34]. In addition to the onsite Hubbard repulsion, our model includes the next-nearest-neighbor (NNN) attraction, which can arise effectively through electron-phonon coupling [35, 36, 37, 38] or fluctuations of the neighboring insulating phase [39, 40, 41]. We then perform a thorough mean-field analysis to the interacting model and obtain the phase diagram at filling $\nu=-1$ under realistic interaction strengths, concluding that the SC phase is consistent with inter-valley pairing with mixed $p_{x}\pm ip_{y}$ and $d_{xy}\mp id_{x^{2}-y^{2}}$-wave symmetry, featuring non-trivial topology under Altland-Zirnbauer A-class with Chern number $C=\pm 2$, and the correlated insulator phase has $120^{\circ}$ antiferromagnetic (AFM) order. Additionally, the mixed symmetry character of the SC phase near zero displacement field predicted by our model has experiment observable distinctions from that of moiré Hubbard model [42, 43, 16, 17, 18] due to the absence of emergent spin-valley SU(2) symmetry [26, 27], providing new insights into the nature of the moiré TMD systems.
Continuum model and symmetry analysis.—We begin with the standard low-energy continuum model description of tWSe2 [6]. The non-interacting continuum Hamiltonian for $K$ valley is given by:
$H_{K}(\bm{r})=\left(\begin{array}[]{cc}-\frac{\hbar^{2}\left(\bm{k}-\bm{\kappa}_{+}\right)^{2}}{2m^{*}}+\Delta_{+}(\bm{r})&\Delta_{\text{T}}(\bm{r})\\ \Delta_{\text{T}}^{\dagger}(\bm{r})&-\frac{\hbar^{2}\left(\bm{k}-\bm{\kappa}_{-}\right)^{2}}{2m^{*}}+\Delta_{-}(\bm{r})\end{array}\right)$ (1)
where $m^{*}$ is the effective mass of valence band edge, $\bm{\kappa}_{\pm}$ are located at corners of the mini Brillouin zone as shown in Fig. 1(a). In the lowest-order harmonic approximation, we only keep terms with moiré wave vectors $\bm{g}_{j}$ which are obtained by rotation of $\bm{g}_{1}=(\frac{4\pi}{\sqrt{3}a_{M}},0)$ by $(j-1)\pi/3$ in moiré potential $\Delta_{\pm}(\bm{r})$ and interlayer tunneling $\Delta_{\text{T}}(\bm{r})$:
$\Delta_{\pm}(\bm{r})$ =±$=\pm\frac{V_{z}}{2}+2V\sum_{j=1,3,5}\cos\left(\bm{g}_{j}\cdot\bm{r}\pm\psi\right)$ (2)
$\Delta_{\text{T}}(\bm{r})$ $=w\left(1+e^{-i\bm{g}_{2}\cdot\bm{r}}+e^{-i\bm{g}_{3}\cdot\bm{r}}\right)$ (3)
and the Hamiltonian for $-K$ valley is related by time-reversal symmetry $\mathcal{T}$. Throughout this paper, we adopt the experimental relevant twist angle $\theta=3.65^{\circ}$ [24], continuum model parameters derived by large scale ab initio simulations ($V$, $\psi$, $w$) = ($9$ meV, $128^{\circ}$, $18$ meV) [29] and the effective mass $m^{*}=0.45m_{e}$ [44]. The moiré lattice constant is given by $a_{M}\approx a_{0}/\theta$ with $a_{0}=3.317$Å [45].
In the absence of displacement field $V_{z}=0$, tWSe2 system has $D_{3}$ point group symmetry generated by threefold rotation $C_{3z}$ and twofold rotation $C_{2y}$. Additionally, due to the lowest order harmonic approximation we adopted, the continuum model has additional pseudo-inversion symmetry $\mathcal{I}$ with $\sigma_{x}H_{K}(\bm{r})\sigma_{x}=H_{K}(-\bm{r})$, enlarging the point group symmetry to $D_{3d}$. When a finite displacement field $V_{z}\neq 0$ is applied, the point group symmetry is reduced from $D_{3d}$ to $C_{3v}$, breaking all symmetries that interchange the two layers. Apart from point group symmetries, the system also exhibits U(1) spin-valley symmetry and time reversal symmetry $\mathcal{T}$ for both $\mathcal{V}_{z}=0$ and $\mathcal{V}_{z}\neq 0$. These symmetry considerations are crucial for the subsequent construction of tight binding model and the classification of SC pairing symmetries.
Wannier functions and tight-binding model.—To study the low-energy physics, we first focus on the top moiré valence bands of the tWSe2 continuum model at $V_{z}=0$. The Chern numbers for each valley share the same sign up to top five moiré bands, precluding any low-energy real-space description with correct band topology due to Wannier obstructions. Therefore, we focus on accurately reproducing the Chern numbers of the top two moiré bands, and construct a minimal three-band tight-binding model with Chern numbers ($1,1,-2$) for $K$ valley. Following Refs. [29, 30, 31], we construct sufficiently localized Wannier functions using their layer polarization properties and the location of Wannier centers. The resulting Wannier functions are shown in Fig. 1(b), where the A and B orbitals are centered at XM and MX regions (±$(\pm a_{M}/\sqrt{3},0)$ with opposite layer polarization, and the (more extended) C orbital is centered at MM region $(0,0)$ with layer hybridization. Detailed symmetry analysis [31] reveals that the (A, B, C) orbitals for the $K$ valley have $C_{3z}$ eigenvalues ($e^{-i2\pi/3}$, $e^{-i2\pi/3}$, $1$), and $C_{2y}\mathcal{T}$ or $\mathcal{I}$ symmetries interchange A and B orbitals but leave C orbital invariant. Using these Wannier functions, we can construct the following tight-binding model:
$H_{0}=\sum_{i\alpha j\beta\sigma}t_{i\alpha j\beta\sigma}c^{\dagger}_{i\alpha\sigma}c_{j\beta\sigma}$ (4)
where $i,j$ are unit cell indexes, $\alpha,\beta$ are sublattice indexes (A, B, C), and $\sigma$ is spin-valley index $\uparrow$ (or $K$), $\downarrow$ (or $-K$). By keeping up to fifth-nearest-neighbor hopping parameters [46], Fig. 1(c) illustrates the close matching between the band structure of the tight-binding model and the continuum model, especially for the top moiré valence band.
Figure 1: (a) Mini Brillouin zone formed by a small twist angle $\theta$ between two layers. (b) Density distribution of Wannier functions on two layers, with unit cell shown as dashed line. (c) Comparison of the band structure between continuum model (dashed line) and tight binding model (solid line), where colors representing sublattice components.
We then consider the case of $V_{z}\neq 0$. Instead of repeating the above procedure for each $V_{z}$ separately, we utilize the layer polarization properties of Wannier functions, modeling the effect of displacement field as:
$H_{D}=\frac{\mathcal{V}_{z}}{2}\sum_{i}(n_{iA}-n_{iB})$ (5)
where $\mathcal{V}_{z}$ is the energy difference between A and B sublattices induced by the displacement field, and $n_{i\alpha}=\sum_{\sigma}c_{i\alpha\sigma}^{\dagger}c_{i\alpha\sigma}$ is the density operator. Since the energy expectation value of displacement field on the C sublattice can be chosen as 00, we neglect terms involving it in $H_{D}$ [47]. The Fermi surfaces of filling factor $\nu=-1$ with $\mathcal{V}_{z}=0$, $5$, $15$ and $25$ meV are shown in Fig. 2(a), where the Fermi surface of spin up and down are split by the displacement field $\mathcal{V}_{z}$, but related by the time reversal symmetry $\mathcal{T}$. Fig. 2(b) also illustrates the Fermi surface density of states (DOS) as a function of $\mathcal{V}_{z}$.
Figure 2: (a) Fermi surfaces of free Hamiltonian at different displacement fields, with colors indicating the sublattice components, and thickness representing the DOS. The spin of the Fermi surfaces are labeled as solid arrows. When $\mathcal{V}_{z}=0$ meV, where spin up and down Fermi surfaces coincide, only spin down Fermi surface is shown. The approximate nesting wave vector $\bm{Q}$ between spin up and down Fermi surfaces are illustrated as dashed arrow. (b) The Fermi surface DOS of A, B, C sublattices as a function of displacement field $\mathcal{V}_{z}$.
To capture the many-body physics, we consider leading interactions in tWSe2. The dominant interaction should be the onsite Hubbard repulsion:
$H_{U}=\sum_{i\alpha}U_{\alpha}n_{i\alpha\uparrow}n_{i\alpha\downarrow}$ (6)
where $\alpha=A,B,C$ labels the sublattice, and $n_{i\alpha\sigma}=c_{i\alpha\sigma}^{\dagger}c_{i\alpha\sigma}$. We take $U_{A}=U_{B}>U_{C}$ due to the Wannier orbital distributions as discussed previously. The positive Hubbard interaction can typically lead to magnetic ordering, providing a promising explanation of the correlated insulator phase.
Due to the time reversal symmetry $\mathcal{T}$ of the systems, the Fermi surface features Cooper instability, namely SC instabilities can occur even with infinitesimal attractions, which can be effectively induced by electron-boson (phonon, magnon, etc.) couplings. Considering that the Fermi surface shown in Fig. 2(a) favor pairing within the same sublattices [46] and has almost no C-orbital component, a natural choice is to consider the attractions on NNN sites of the same A or B sublattices:
$H_{V}=-V\sum_{i}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in\text{NNN}}n_{i+\delta\alpha}n_{i\alpha}$ (7)
where $V$ is the strength of NNN attraction, $\alpha$ is sublattice index of A and B, and $\delta$ represents one of the three NNN bond directions $120^{\circ}$ apart. Besides, we have also considered the effects of nearest-neighbor (NN) attraction between A and B sublattices in the Supplemental Material [46], where we have shown that it is less dominant. The interacting model which we should consider to describe the main physics of tWSe2 is denoted as $H=H_{0}+H_{D}+H_{U}+H_{V}.$
Figure 3: (a) The NNN SC order parameters $\Delta_{AA}$ and $\Delta_{BB}$, where the A, B, C sites are labeled as red, green, blue dots. (b) The pairing form factor of $s$-wave, $f$-wave, $p+ip$-wave, and $d-id$-wave on NNN bonds, with the irreducible representation labeled. (c) $120^{\circ}$ AFM pattern on a certain type of sublattice with wave vector ±$\pm Q$. (d) The folded Brillouin zone induced by the AFM order.
Mean-field analysis of SC and AFM— We start with the mean-field analysis of SC instabilities by decoupling the NNN attraction $H_{V}$ in the SC channel as:
$H_{V}\approx-V\sum_{i\sigma\sigma^{\prime}}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in NNN}\bigg{(}\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}^{*}(\delta)c_{i\alpha\sigma^{\prime}}c_{i+\delta\alpha\sigma}$
$~{}~{}~{}~{}~{}~{}+c_{i+\delta\alpha\sigma}^{\dagger}c_{i\alpha\sigma^{\prime}}^{\dagger}\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)-\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}^{*}(\delta)\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)\bigg{)},$ (8)
where $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)=\langle c_{i\alpha\sigma^{\prime}}c_{i+\delta\alpha\sigma}\rangle$ is the spatially uniform real-space pairing order parameter on NNN bond $\delta$, as shown in Fig. 3(a). To further determine the ansatz of the pairing order parameter $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)$, we analyze the possible pairing symmetries, which should be classified by the irreducible representations of the symmetry group. Since the addition of displacement field $\mathcal{V}_{z}$ preserves time reversal symmetry $\mathcal{T}$ but breaks pseudo-inversion symmetry $\mathcal{I}$, we only consider $S_{z}=0$ sector for the U(1) spin-valley symmetry (i.e. inter-valley pairing). And these $S_{z}=0$ pairings can be further categorized into singlet pairing $\Delta_{\alpha}^{S}(\delta)=(\tilde{\Delta}_{\alpha\uparrow\downarrow}(\delta)-\tilde{\Delta}_{\alpha\downarrow\uparrow}(\delta))/\sqrt{2}$ and triplet pairing $\Delta_{\alpha}^{T}(\delta)=(\tilde{\Delta}_{\alpha\uparrow\downarrow}(\delta)+\tilde{\Delta}_{\alpha\downarrow\uparrow}(\delta))/\sqrt{2}$ [48]. For the point group symmetry $C_{3v}$ under finite displacement field, a classification of pairing symmetries based on momentum space continuum model has been provided in Ref. [49]. Switching to our tight-binding description, it is straightforward to show that only $A_{1}$ (mixing $s$ and $f$-wave) and $E$ (mixing $(p_{x},p_{y})$ and $(d_{xy},d_{x^{2}-y^{2}})$-wave) representations are possible for NNN pairing. The SC pairing form factors, focusing only on chiral SC in $E$ representation, are illustrated in Fig. 3(b), and we leave the discussion of nematic SC in $E$ representation in the Supplemental Material [46], where we have shown it is subdominant.
To understand the competing insulating phase, we decouple the onsite Hubbard repulsion $H_{U}$ in the channel of magnetic ordering in the $xy$-plane [26, 27] as:
$H_{U}=-\sum_{i\alpha}U_{\alpha}\left(m_{i\alpha}c_{i\alpha\downarrow}^{\dagger}c_{i\alpha\uparrow}+m_{i\alpha}^{*}c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}\right)$ (9)
$~{}~{}~{}~{}~{}~{}~{}~{}~{}+\sum_{i\alpha}U_{\alpha}|m_{i\alpha}|^{2}+\frac{1}{2}\sum_{i\alpha\sigma}U_{\alpha}n_{i\alpha\sigma}$
where $m_{i\alpha}=\langle c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}\rangle=\langle S_{i\alpha}^{x}\rangle+i\langle S_{i\alpha}^{y}\rangle$ is the complex order parameter representing the in-plane magnetization. We further constraint the form of $m_{i\alpha}$ by noting that, as shown in Fig. 2(a), although the Fermi surface deforms as the displacement field $\mathcal{V}_{z}$ changes, an approximate nesting between spin up and spin down Fermi surfaces persist. Moreover, the nesting wave vectors are relatively close to the commensurate ones ±$\pm\bm{Q}=(0,\pm 4\pi/3a_{M})$ over a wide range of displacement field $\mathcal{V}_{z}$, giving raise to the $120^{\circ}$ AFM order $m_{i\alpha}=m_{\alpha}^{+}e^{iQ\cdot R_{i}}+m_{\alpha}^{-}e^{-iQ\cdot R_{i}}$ with two possible chiralities [50] as shown in Fig. 3(c). Such AFM order breaks translation symmetry of the system, folding the Brillouin zone as shown in Fig. 3(d), with the new filling factor in terms of the folded Brillouin zone becoming $\tilde{\nu}=-3$. Depending on details of the system, the ground state can be either metallic or insulating [51] due to the nesting of ±$\pm\bm{Q}$ is non-perfect.
We then perform mean-field calculations for SC and AFM independently, using intermediate Hubbard repulsion $U_{A}=U_{B}=35$ meV, $U_{C}=25$ meV and two different NNN attractions $V=10$ meV or $V=12.5$ meV as representing parameters. The resulting phase diagrams for tuning displacement field $\mathcal{V}_{z}$ at filling $\nu=-1$ are summarized in Fig. 4(e) and (f), respectively. And we leave more detailed mean-field derivations and discussions in the Supplemental Material [46].
Figure 4: (a) The energy gain per hole of the ordered phases compared to the symmetric phase. (b) SC orders for $V_{2}=10$ meV on one specific NNN bond $\delta_{0}$ under the gauge described in the main text. The sign of order parameters correspond to the sign of real(imaginary) part of singlet(triplet) pairings. (c) The AFM mean-field charge gap at filling factor $\tilde{\nu}=-3$ in the folded Brillouin zone. (d) Magnitude of the AFM orders on A, B, C sublattices. (e, f) The phase diagrams for $V_{2}=10$ meV and $V_{2}=12.5$ meV respectively, indicating that under mean-field framework, the SC to AFM-I transition can either have an intermediate AFM-M phase or occur directly.
Topological superconductivity.—We first focus on the SC phase in zero or small displacement field regime. A comparison of the energy gain $\Delta E$ between possible SC orders and $120^{\circ}$ AFM order is shown in Fig. 4(a), which indicates that the system is in SC phase for $\mathcal{V}_{z}<\mathcal{V}_{z,c}\approx 4.0$ meV ($5.2$ meV) for $V=10$ meV ($12.5$ meV), qualitatively consistent with the critical field $\mathcal{V}_{z,c}^{\text{exp}}\approx 2.6$ meV observed in the experiment [24]. The pairing of this SC phase is chiral with mixed $p_{x}\pm ip_{y}$ and $d_{xy}\mp id_{x^{2}-y^{2}}$-wave in the $E$ representation of $C_{3v}$, spontaneously breaking the time reversal symmetry $\mathcal{T}$. More importantly, such chiral SC phase, which fits into the A-class (namely, breaking $\mathcal{T}$ but preserving $S_{z}$) of the Altland-Zirnbauer tenfold classification scheme [52, 53], is topologically non-trivial with Chern number computed to be $C=\pm 2$, suggesting tWSe2 as a promising candidate for realizing chiral topological SC. The internal structure of the SC order as a function of $\mathcal{V}_{z}$ for $V=10$ meV is also illustrated in Fig. 4(b), where the consistency constraint between time reversal symmetry $\mathcal{T}$ and $C_{3v}$ mirror plane symmetry enables us to fix a simple gauge such that all singlet pairings $\Delta^{S}_{\alpha}(\delta)$ are real while all triplet pairings $\Delta^{T}_{\alpha}(\delta)$ are imaginary. It is worth emphasizing that, unlike the moiré Hubbard model [42, 43, 16, 17, 18], the mixing of singlet and triplet pairings are allowed for all $\mathcal{V}_{z}$, especially for $\mathcal{V}_{z}=0$ due to the absence of emergent spin-valley SU(2) symmetry [54], and the dominance of $d_{xy}\mp id_{x^{2}-y^{2}}$ component over the $p_{x}\pm ip_{y}$ one in the SC phase is in accordance with the result of Chern number ±$\pm 2$. Moreover, the SC is enhanced (suppressed) on the B (A) sublattice upon increasing the displacement field $\mathcal{V}_{z}$ from zero, closely following the trend of Fermi surface DOS in Fig. 2(b).
Antiferromagnetic correlated insulators.—As the displacement field is further increased beyond a critical field $\mathcal{V}_{z,c}$, the system transitions into the $120^{\circ}$ AFM phase. To further determine its transport properties, we compute the AFM mean-field charge gap for $\tilde{\nu}=-3$. As shown in Fig. 4(c), a broad intermediate range of AFM insulator (AFM-I) phase appears for $\mathcal{V}_{z,c}\lesssim\mathcal{V}_{z}<\mathcal{V}^{\prime}_{z,c}\approx 20.8$ meV and an AFM metal (AFM-M) phase for $\mathcal{V}_{z}>\mathcal{V}^{\prime}_{z,c}$, closely matching the phenomenology of the experiment [24]. In the mean-field framework, depending on the value of $V$ employed in the model, the SC to AFM-I transition either exhibits a tiny intermediate AFM-M phase as shown in Fig. 4(e) or occurs as a direct first-order transition as shown in Fig. 4(f). To explain the transport evidences of continuous superconductor-insulator transition [24], disorder could play an important role. The continuous transition into the SC phase might potentially be a percolation transition [55, 56] of local SC regions induced by disorders or a disordered-round first-order transition. We also present the magnitude of AFM orders on different sublattices as a function of $\mathcal{V}_{z}$ in Fig. 4(d). Except for small $\mathcal{V}_{z}$, where the AFM order is stronger on B sublattice due to its larger DOS (see Fig. 2(b)), the AFM order generally favors A sublattice, as the holes are mostly concentrated there in response to the displacement field $\mathcal{V}_{z}$, as well as its better approximation for commensurate nesting wave vector $\bm{Q}$. And the sudden drop of AFM orders at large $\mathcal{V}_{z}$ regime coincides with the disappearance of B sublattice hole pockets at $\kappa_{\pm}$ points and the sharp decline in B sublattice DOS, as illustrated in Fig. 2, indicating that the holes on B sublattice play a crucial role in mediating the AFM order.
Summary and Discussions.— In summary, we have constructed a three-band tight-binding model through direct Wannierization, and incorporated onsite Hubbard repulsion and NNN attraction to explain the SC and correlated insulator phase observed in the $3.65^{\circ}$ tWSe2 at filling $\nu=-1$ [24]. Our mean-field analysis indicates that, the SC phase is an A-class topological SC with Chern number $C=\pm 2$, featuring the inter-valley mixed $p_{x}\pm ip_{y}$ and $d_{xy}\mp id_{x^{2}-y^{2}}$-wave pairing symmetry, and the correlated insulator phase is explained by the $120^{\circ}$ AFM order. The topological band structure in our model can be crucial for understanding how SC arises for flat band systems, since non-trivial lower bounds of SC superfluid weight exists [57, 58, 59] due to quantum geometric effects, which deserves more detailed future theoretical studies. Moreover, our results also suggest that the tWSe2 homobilayers could provide new possibilities for realizing topological SC, which call for more detailed experimental studies to further uncover its nature. Identifying the topological edge states by measuring quantized thermal Hall conductance [60], or verifying the chiral nature of the SC pairing symmetry through phase sensitive Josephson junction [61], will surely open new opportunities in twisted TMD systems.
Note added.—In finishing the present work, we noticed that Refs. [62, 63, 49, 64, 65, 66] appeared, which also investigated the SC and correlated insulators in tWSe2 reported in Ref. [24], although there are important differences between those studies and the present one.
Acknowledgements—We would like to thank Andrei Bernevig for helpful discussions. This work is supported in part by the Innovation Program for Quantum Science and Technology (grant No. 2021ZD0302502), by MOSTC under Grant No. 2021YFA1400100, by NSFC under Grant Nos. 12347107 and 12334003. HY acknowledgement the support in part by the Xplorer Prize through the New Cornerstone Science Foundation.
Supplemental Material for ‘Theory of Topological Superconductivity and Antiferromagnetic Correlated Insulators in Twisted Bilayer WSe2
I Hopping Parameters of the Tight-Binding Model
In this section, we provide additional information of the tight-binding model we adopted in the main text, derived from direct Wannierization of the continuum model for $3.65^{\circ}$ tWSe2 at $V_{z}=0$. Fig. S1 illustrates the symmetry inequivalent representative hopping bonds up to $5$th-nearest-neighbor, and we provide the hopping parameters for $K$ valley on these representative bonds as well as the onsite potentials as:
$t_{0}^{A}=t_{0}^{B}=-15.26\text{ meV},\quad t_{0}^{C}=-42.89\text{ meV}$ (S1)
$t^{BC}_{1}=7.12\text{ meV},\quad t^{CC}_{\sqrt{3}}=2.76\text{ meV},\quad t^{AC}_{2}=1.15\text{ meV}$
$t^{AC}_{\sqrt{7}}=3.01e^{i0.965\pi}\text{ meV},\quad t^{CC}_{3}=0.43\text{ meV}$
$t^{AB}_{1}=-4.99\text{ meV},\quad t^{BB}_{\sqrt{3}}=5.16e^{i0.618\pi}\text{ meV}$
$t^{AB}_{2}=0.7\text{ meV},\quad t^{AB}_{\sqrt{7}}=0.78\text{ meV},\quad t^{BB}_{3}=-0.31\text{ meV}$
where $t_{d}^{\alpha\beta}$ represents the hopping parameter from $\beta$ sublattice to $\alpha$ sublattice with distance $d$ (set NN with $d=1$), and $t_{0}^{\alpha}$ denotes the onsite potential of $\alpha$ sublattice.
There are several remarks on the symmetry properties of these hopping parameters. First, hopping parameters related by $C_{3z}$ symmetry will (will not) be identical if $\alpha,\beta$ have identical (different) $C_{3z}$ eigenvalues (see main text). Second, $C_{2y}\mathcal{T}$ and $\mathcal{I}$ symmetries constrain most but not all hopping parameters to be real. Third, the hopping parameters for $K$ and $-K$ valley are related by time reversal symmetry $\mathcal{T}$. Last, unlike the single-band moiré Hubbard model, where emergent spin-valley SU(2) symmetry exists at $V_{z}=0$, the spin-valley symmetry here remains U(1) due to the presence of complex hopping parameters.
Figure S1: Illustration of the representative hopping bonds up to $5$th-nearest-neighbor.
II Derivations and Supplemental Results of Superconducting Mean-Field Analysis
As discussed in the main text, we focus on the interacting tight-binding Hamiltonian of tWSe2 with non-interacting term $H_{0}$, displacement field term $H_{D}$, onsite Hubbard repulsion $H_{U}$, and NNN attraction within A and B sublattices $H_{V_{2}}$:
$H=H_{0}+H_{D}+H_{U}+H_{V_{2}}=\sum_{ij\alpha\beta\sigma}\left(t_{i\alpha j\beta\sigma}-\mu_{\alpha}\delta_{ij}\delta_{\alpha\beta}\right)c_{i\alpha\sigma}^{\dagger}c_{j\beta\sigma}+\sum_{i\alpha}U_{\alpha}n_{i\alpha\uparrow}n_{i\alpha\downarrow}-V_{2}\sum_{i\sigma\sigma^{\prime}\alpha}\sum_{\in\{A,B\}}\sum_{\delta\in NNN}n_{i+\delta\alpha\sigma}n_{i\alpha\sigma^{\prime}},$ (S2)
where we have combined the chemical potential and displacement field as $\mu_{A}=\mu-\mathcal{V}_{z}/2$, $\mu_{B}=\mu+\mathcal{V}_{z}/2$ and $\mu_{C}=\mu$ for simplicity, and other notations are the same as the main text.
The NNN attraction $H_{V_{2}}$ generally leads to a SC phase, while the onsite Hubbard repulsion $H_{U}$ typically favors magnetic ordered phase ($120^{\circ}$ AFM order in this case). Since the SC phase and magnetic ordered phase typically do not coexist, here we assume vanishing magnetic order in the SC mean-field analysis. Thus, according to Eq. (9) in the main text, the mean-field decoupling of $H_{U}$ only contributes an additional sublattice potential term $\frac{1}{2}\sum_{i\alpha}U_{\alpha}n_{i\alpha}$.
We decouple the NNN attraction $H_{V_{2}}$ in the SC channel as:
$H_{V_{2}}\approx-V_{2}\sum_{i\sigma\sigma^{\prime}}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in NNN}\left(\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}^{*}(\delta)c_{i\alpha\sigma^{\prime}}c_{i+\delta\alpha\sigma}+c_{i+\delta\alpha\sigma}^{\dagger}c_{i\alpha\sigma^{\prime}}^{\dagger}\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)-\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}^{*}(\delta)\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)\right),$ (S3)
where we have defined the spatially uniform real space pairing order parameter as $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)=\langle c_{i\alpha\sigma^{\prime}}c_{i+\delta\alpha\sigma}\rangle$.
The dependence of order parameter $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)$ on NNN bond direction $\delta$ is classified according to the SC pairing symmetries. In the main text, we have classified the $S_{z}=0$ inter-valley pairing based on the irreducible representation of the point group $C_{3v}$. Additionally, here we also consider the case that the SC phase breaks $C_{3}$ rotation symmetry of the system (i.e. nematic SC) in E representation of $C_{3v}$, under the assumption that the pairing still preserves one of the three mirror symmetry of $C_{3v}$. Tab. S1 summarizes the $C_{3v}$ irreducible representations and the form factors $f(\delta)$ (defined by $\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}(\delta)=f(\delta)\tilde{\Delta}_{\alpha\sigma\sigma^{\prime}}$) for each pairing symmetries we considered. And in the following calculations, we will examine each pairing symmetries separately, comparing their energies to determine the most favorable one.
Pairing Symmetry Irreducible Representation $f(\delta_{1})$ $f(\delta_{2})$ $f(\delta_{3})$
Mixed $s$-$f$ wave $A_{1}$ $1$ $1$ $1$
Mixed $p_{x}\pm ip_{y}$ $d_{xy}\mp id_{x^{2}-y^{2}}$ wave $E$ $1$ $e^{\pm i\frac{2\pi}{3}}$ $e^{\mp i\frac{2\pi}{3}}$
Mixed $p_{x}$ $d_{xy}$ wave $E$ 00 $\frac{\sqrt{3}}{2}$ $-\frac{\sqrt{3}}{2}$
Mixed $p_{y}$ $d_{x^{2}-y^{2}}$ wave $E$ $1$ $-\frac{1}{2}$ $-\frac{1}{2}$
Table S1: Details of SC pairing symmetries on NNN bonds
With the above symmetry considerations, we take $\sigma^{\prime}=\bar{\sigma}$ to be opposite to $\sigma$ for $S_{z}=0$ inter-valley pairing, and then Fourier transform the real space mean-field decoupling Eq. (S3) to momentum space:
$\begin{split}&-V_{2}\sum_{i\sigma}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in NNN}\left(\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}^{*}(\delta)c_{i\alpha\bar{\sigma}}c_{i+\delta\alpha\sigma}+c_{i+\delta\alpha\sigma}^{\dagger}c_{i\alpha\bar{\sigma}}^{\dagger}\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}(\delta)-\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}^{*}(\delta)\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}(\delta)\right)\\ =&\sum_{k\sigma}\sum_{\alpha\in\{A,B\}}\left(\Delta_{\alpha\sigma\bar{\sigma}}^{*}(k)c_{-k\alpha\bar{\sigma}}c_{k\alpha\sigma}+c_{k\alpha\sigma}^{\dagger}c_{-k\alpha\bar{\sigma}}^{\dagger}\Delta_{\alpha\sigma\bar{\sigma}}(k)\right)+NV_{2}\sum_{\sigma}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in NNN}\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}^{*}(\delta)\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}(\delta)\end{split}$ (S4)
where we have defined the momentum space gap function $\Delta_{\alpha\sigma\bar{\sigma}}(k)$ as $\Delta_{\alpha\sigma\bar{\sigma}}(k)\equiv-V_{2}\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}\sum_{\delta\in NNN}f(\delta)e^{-ik\delta}$, and the self-consistency relation in momentum space is given by:
$\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}(\delta)=\frac{1}{N}\sum_{k}\langle c_{-k\alpha\bar{\sigma}}c_{k\alpha\sigma}\rangle e^{ik\delta}.$ (S5)
Together with the Fourier transformation of $H_{0}+H_{D}$, we can then write the full momentum space mean-field Hamiltonian into the Bogoliubov-de-Gennes(BdG) form:
$H_{\text{MF}}=$ $\sum_{k\alpha\beta\sigma}\left(\epsilon_{\alpha\beta\sigma}(k)-\mu_{\alpha}\delta_{\alpha\beta}\right)c_{k\alpha\sigma}^{\dagger}c_{k\beta\sigma}+\sum_{k\sigma}\sum_{\alpha\in\{A,B\}}\left(\Delta_{\alpha\sigma\bar{\sigma}}^{*}(k)c_{-k\alpha\bar{\sigma}}c_{k\alpha\sigma}+c_{k\alpha\sigma}^{\dagger}c_{-k\alpha\bar{\sigma}}^{\dagger}\Delta_{\alpha\sigma\bar{\sigma}}(k)\right)$ (S6)
$+NV_{2}\sum_{\sigma}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in NNN}\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}^{*}(\delta)\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}(\delta)$
$=$ $\frac{1}{2}\sum_{k}\begin{pmatrix}\bm{c}_{k\uparrow}^{\dagger}&\bm{c}_{k\downarrow}^{\dagger}&\bm{c}_{-k\uparrow}&\bm{c}_{-k\downarrow}\end{pmatrix}\begin{pmatrix}\bm{\epsilon}_{\uparrow}(k)-\bm{\mu}&\bm{0}&\bm{0}&\bm{\Delta}_{\uparrow\downarrow}(k)\\ \bm{0}&\bm{\epsilon}_{\downarrow}(k)-\bm{\mu}&\bm{\Delta}_{\downarrow\uparrow}(k)&\bm{0}\\ \bm{0}&\bm{\Delta}_{\downarrow\uparrow}^{\dagger}(k)&-\bm{\epsilon}^{T}_{\uparrow}(-k)+\bm{\mu}&\bm{0}\\ \bm{\Delta}_{\uparrow\downarrow}^{\dagger}(k)&\bm{0}&\bm{0}&-\bm{\epsilon}_{\downarrow}^{T}(-k)+\bm{\mu}\end{pmatrix}\begin{pmatrix}\bm{c}_{k\uparrow}\\ \bm{c}_{k\downarrow}\\ \bm{c}_{-k\uparrow}^{\dagger}\\ \bm{c}_{-k\downarrow}^{\dagger}\end{pmatrix}$
$+\frac{1}{2}\sum_{k\alpha\sigma}\left(\epsilon_{\alpha\alpha\sigma}(k)-\mu_{\alpha}\right)+NV_{2}\sum_{\sigma}\sum_{\alpha\in\{A,B\}}\sum_{\delta\in NNN}\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}^{*}(\delta)\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}(\delta),$
where the pairing matrix is defined as:
$\bm{\Delta}_{\sigma\bar{\sigma}}(k)=\begin{pmatrix}\Delta_{A\sigma\bar{\sigma}}(k)-\Delta_{A\bar{\sigma}\sigma}(-k)&0&0\\ 0&\Delta_{B\sigma\bar{\sigma}}(k)-\Delta_{B\bar{\sigma}\sigma}(-k)&0\\ 0&0&0\end{pmatrix}$ (S7)
With the mean-field Hamiltonian derived above, we can perform the standard mean-field calculations. We first construct the BdG Hamiltonian in Eq. (S6) using some randomly chosen initial values of the order parameters $\tilde{\Delta}_{\alpha\sigma\bar{\sigma}}$, and adjust the chemical potential $\mu$ to fit the correct average filling $\nu=-1$ by $n_{e}=\sum_{k\alpha\sigma}\langle c_{k\alpha\sigma}^{\dagger}c_{k\alpha\sigma}\rangle$, with the many-body ground state of the BdG Hamiltonian is obtained by filling all quasi-particle states with negative energy. And the ground state expectation value in Eq. (S5) is calculated to update the order parameters for the next iteration. The above process is repeated until the self-consistency is achieved.
Although we have argued in the main text that NNN attraction is more likely to induce stronger SC instability, here we also examine an alternative scenario that the NNN attraction $H_{V_{2}}$ is replaced by a NN attraction $H_{V_{1}}$:
$H^{\prime}=H_{0}+H_{D}+H_{U}+H_{V_{1}}=\sum_{ij\alpha\beta\sigma}\left(t_{i\alpha j\beta\sigma}-\mu_{\alpha}\delta_{ij}\delta_{\alpha\beta}\right)c_{i\alpha\sigma}^{\dagger}c_{j\beta\sigma}+\sum_{i\alpha}U_{\alpha}n_{i\alpha\uparrow}n_{i\alpha\downarrow}-V_{1}\sum_{i\sigma\sigma^{\prime}}\sum_{\delta\in NN}n_{i+\delta A\sigma}n_{iB\sigma^{\prime}}.$ (S8)
The SC mean-field analysis for $H^{\prime}$ closely mirrors that of $H$ we have derived above, where the only differences are the substitution of $\delta$ from NNN to NN bonds and the modification of sublattice indices. Therefore, we will omit the derivation here for brevity, and only show the corresponding classification of SC pairing symmetries on NN bonds (only considering chiral SC in E representation) in Tab. S2.
Pairing Symmetry Irreducible Representation $f(\delta_{1})$ $f(\delta_{2})$ $f(\delta_{3})$
$s$ wave $A_{1}$ $1$ $1$ $1$
Mixed $p_{x}\pm ip_{y}$ $d_{xy}\mp id_{x^{2}-y^{2}}$ wave $E$ $1$ $e^{\pm i\frac{2\pi}{3}}$ $e^{\mp i\frac{2\pi}{3}}$
Table S2: Details of SC pairing symmetries on NN bonds
Taking $U_{A}=U_{B}=35$ meV, $U_{C}=25$ meV, and $V_{2}=10$ meV as in the main text for the NNN Hamiltonian $H$ (see Eq. (S2)), and $V_{1}=10$ meV for the NN Hamiltonian $H^{\prime}$ (see Eq. (S8)), we perform the mean-field calculations for all pairing symmetries classified in Tab. S1 and Tab. S2. The resulting energy gain per hole for NNN and NN Hamiltonians are illustrated in Fig. S2(a) and (b) respectively.
A direct comparison of the energy gains $\Delta E$ reveals that NNN pairings are approximately $2\sim 3$ orders of magnitude stronger than NN pairings for same attraction strength $V_{1}=V_{2}$. We thus conclude that the NNN pairings should be more dominant than the NN pairings, justifying our main focus of the NNN pairings in the main text. This result is actually straightforward to understand. Due to the time reversal symmetry $\mathcal{T}$ in both $H$ and $H^{\prime}$, the Cooper pairs will form between time-reversal-related states. As the time reversal symmetry $\mathcal{T}$ does not alter sublattice index $\alpha$, and given that the Fermi surface have relatively strong sublattice polarization property (as shown in Fig. 2(a) in the main text), the system is expected to favor intra-sublattice pairings (e.g. NNN pairings) over inter-sublattice pairings (e.g. NN pairings). Considering the layer polarization property of the A and B sublattices (the main components of the Fermi surface), it is equivalent to say that the tWSe2 system favors intra-layer pairings.
Next, we examine the nematic SC orders (with mixed $p_{x}$-$d_{xy}$ or $p_{y}$-$d_{x^{2}-y^{2}}$-wave symmetry) as shown in Fig. S2(a). These nematic SC states have energies considerably higher than the chiral ones under small displacement field. It is expected that some moderate $C_{3}$ breaking terms (e.g. strain) may favor such nematic SC over chiral ones. And it should be also noted that these two nematic SC states are nearly (but not exactly) degenerate in energy.
Returning to the non-nematic cases for NNN pairings, one can see that at small $\mathcal{V}_{z}$ (as discussed in the main text) or large enough $\mathcal{V}_{z}$ regime, the chiral mixed $p_{x}\pm ip_{y}$ $d_{xy}\mp id_{x^{2}-y^{2}}$-wave symmetry is the most dominant, while at intermediate $\mathcal{V}_{z}$, the mixed $s$-$f$-wave symmetry becomes the strongest one. To understand this, we plot the magnitude of the momentum space form factor for NNN pairings with different partial wave components in Fig. S3. As shown in Fig. S3, the $s$-wave component has local maximum at $\gamma$ and $\kappa_{\pm}$ points of Brillouin zone(BZ) while having ring-shaped minimum in between; the $p_{x}\pm ip_{y}$-wave component is large except at $\gamma$ points and the BZ boundary; the $d_{xy}\pm id_{x^{2}-y^{2}}$-wave component has maximum at $m$ points while having minimum at $\gamma$ and $\kappa_{\pm}$ points; and the $f$ wave component has maximum at $\kappa_{\pm}$ points but having minimum nodal line along $\gamma$-$m$ directions.
Figure S2: The energy gain $\Delta E$ per hole of the SC phases compared to the symmetric phase with (a) NNN and (b) NN pairings.
Figure S3: Momentum space form factor for NNN pairing with (a) s-wave, (b) $p_{x}\pm ip_{y}$-wave, (c) $d_{xy}\mp id_{x^{2}-y^{2}}$-wave and (d) $f$-wave symmetry, and the BZ is shown as black dash line in each figure.
The energetic of these different SCs can then be understood by combining the Fermi surface structures shown in Fig. 2 in the main text with the form factors shown in Fig. S3. At small displacement field $\mathcal{V}_{z}$, the Fermi surface is hexagonal shape around the $\gamma$ point in the BZ, where the chiral $p_{x}\pm ip_{y}$ and $d_{xy}\pm id_{x^{2}-y^{2}}$ components are the largest and fully gaps the Fermi surface. When increasing $\mathcal{V}_{z}$, Fig. 2(b) in the main text indicates that B sublattice has dominant DOS, which should be our primary focus. And Fig. 2(a) in the main text clearly illustrates that the B pockets are move towards $\kappa_{\pm}$ points in the BZ as the displacement field $\mathcal{V}_{z}$ is increasing, where the $s$ and $f$-wave components are larger in magnitude and become better in energy. However, as $\mathcal{V}_{z}$ increases further, the B sublattice Fermi pockets at $\kappa_{\pm}$ points disappear, leaving the remaining Fermi surface mainly on A sublattice, which is again located at the region where $p_{x}\pm ip_{y}$ and $d_{xy}\pm id_{x^{2}-y^{2}}$ components are dominant. The above analysis suggests that the combined use of the structure of Fermi surfaces and momentum space pairing form factors provides a clear explanation of the mean-field behaviors of the SC phases.
III Derivations of Magnetic Mean-Field Analysis
In this section, we provide a more detailed mean-field analysis for possible magnetic orders of the NNN Hamiltonian $H$ in Eq. (S2). Assuming the absence of SC order (hence $H_{V_{2}}$ has vanishing mean-field contributions), we mainly focus on the in-plane magnetic channel of $H_{U}$ as:
$H_{U}\approx-\sum_{i\alpha}U_{\alpha}\left(m_{i\alpha}c_{i\alpha\downarrow}^{\dagger}c_{i\alpha\uparrow}+m_{i\alpha}^{*}c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}\right)+\sum_{i\alpha}U_{\alpha}|m_{i\alpha}|^{2}+\frac{1}{2}\sum_{i\alpha\sigma}U_{\alpha}n_{i\alpha\sigma},$ (S9)
where $m_{i\alpha}=\langle c_{i\alpha\uparrow}^{\dagger}c_{i\alpha\downarrow}\rangle=\langle S_{i\alpha}^{x}\rangle+i\langle S_{i\alpha}^{y}\rangle$ is the in-plane magnetic order parameter. As discussed in the main text, due to the existence of approximate nesting wave vector $\bm{Q}=(0,4\pi/3a_{M})$ between spin up and down Fermi surfaces, we first consider the ansatz of $120^{\circ}$ AFM order $m_{i\alpha}=m_{\alpha}^{+}e^{iQ\cdot R_{i}}+m_{\alpha}^{-}e^{-iQ\cdot R_{i}}=\sum_{\eta}m_{\alpha}^{\eta}e^{i\eta Q\cdot R_{i}}$. Performing Fourier transformation, the mean-field Hamiltonian is given by:
$H_{\text{MF}}=\sum_{k\alpha\beta\sigma}\left(\epsilon_{\alpha\beta\sigma}(k)+\frac{U_{\alpha}}{2}\delta_{\alpha\beta}-\mu_{\alpha}\delta_{\alpha\beta}\right)c_{k\alpha\sigma}^{\dagger}c_{k\beta\sigma}-\sum_{k\alpha\eta}U_{\alpha}\left(m_{\alpha}^{\eta}c_{k+\eta Q\alpha\downarrow}^{\dagger}c_{k\alpha\uparrow}+m_{\alpha}^{\eta*}c_{k-\eta Q\alpha\uparrow}^{\dagger}c_{k\alpha\downarrow}\right)+N\sum_{\alpha\eta}U_{\alpha}|m_{\alpha}^{\eta}|^{2},$ (S10)
and the order parameter $m_{\alpha}^{\eta}$ satisfies the self-consistent equation:
$m_{\alpha}^{\eta}=\frac{1}{N}\sum_{k}\langle c_{k\alpha\uparrow}^{\dagger}c_{k+\eta Q\alpha\downarrow}\rangle.$ (S11)
Notice that the momentum summation above should be performed in the original BZ with $N$ being the number of original unit cells. And in practice, the folding of BZ, as shown in Fig. 3(d) in the main text, can be easily achieved by a simple relabeling of momentum.
Then, we can perform the standard mean-field calculations as follow. We first construct and diagonalize the mean-field Hamiltonian in Eq. (S10), and filled the lowest energy states until $\nu=-1$. We then use Eq. S11 to update the order parameters for next iteration. Such process is repeated until self-consistency is achieved, and the numerical results are detailed in the main text.
However, there are also other possible competing orders, one important class being the zero momentum magnetic orders, which do not break the translation symmetry. Since there is no experimental evidence of valley polarization ferromagnetism (i.e. magnetic order in $z$ direction), here we again focus on the in-plane magnetic channel. The decoupling of the Hubbard interaction $H_{U}$ in momentum space is:
$H_{U}\approx-\sum_{k\alpha}U_{\alpha}\left(m_{\alpha}c_{\alpha\vec{k}\downarrow}^{\dagger}c_{\alpha\vec{k}\uparrow}+m_{\alpha}^{*}c_{\alpha\vec{k}\uparrow}^{\dagger}c_{\alpha\vec{k}\downarrow}\right)+N\sum_{\alpha}U_{\alpha}m_{\alpha}^{2}+\frac{1}{2}\sum_{k\alpha\sigma}U_{\alpha}n_{k\alpha\sigma},$ (S12)
where the in-plane magnetic order parameter is given by $m_{\alpha}=\frac{1}{N}\sum_{k}\langle c_{k\alpha\uparrow}^{\dagger}c_{k\alpha\downarrow}\rangle$. We have numerically checked that such magnetic order is not favored when turning the displacement field on, and we conclude that $120^{\circ}$ AFM order is indeed more favorable. Such results can be easily understood since the displacement field splits the Fermi surfaces of spin up and spin down, hence suppress the zero-momentum inter-valley coherent magnetic orders.